# On the Nature of the Fundamental Plasma Excitation in a Plasmonic Crystal

### v1

@@ Abstract
We report on the experimental study of the spectrum of plasma excitations in a plasmonic crystal fabricated from the two-dimensional electron system in an AlGaAs/GaAs semiconductor heterostructure. Our results reveal that the plasmonic crystal supports only a single fundamental plasma mode, challenging the current theoretical understanding that anticipates multiple modes. This unexpected finding is supported by comprehensive research on the mode frequency and relaxation as a function of the gate width across different plasmonic crystal periods. Furthermore, we develop an analytical approach that accurately describes the behavior of plasma excitations in plasmonic crystals, providing new insights into the fundamental physics of plasmonic systems.
@@
The currently prevailing theoretical model suggests that an incident electromagnetic wave excites two types of standing plasma waves in the plasmonic crystal. When the distance from the gate to the 2DES is relatively small, this model is analogous to the tight-binding limit of the Kronig-Penney model, which describes the particle behavior in a periodic potential [21, 22, 23, 24]. In this limit, the fundamental plasma excitations of the plasmonic crystal decouple into two sets of solutions. The first mode is an ordinary unscreened plasmon, which is excited in the gap between the gates, with its frequency described by [25]
where is the 2D electron density, is the effective electron mass, ( is an integer number, is the width of the gate strips) is the plasmon wave vector, is the effective dielectric permittivity of the medium surrounding the 2DES, and is the period of the plasmonic crystal. The second mode is the screened plasmon confined to the region directly underneath the gate, with its spectrum taking the linear form [26]
Here is the distance between the gate and the 2DES, ( is an integer), and is the effective dielectric permittivity of the semiconductor crystal, specifically . The most important conclusion of the Kronig-Penney model is that in the limit of the screened and unscreened plasmon modes can be excited independently in 2DES [21, 22, 23, 24].
In the present work, we investigate the terahertz (THz) spectrum and relaxation of plasma excitations in a plasmonic crystal fabricated from an AlGaAs/GaAs heterostructure. Our experimental data reveal that the plasmonic crystal supports only a single fundamental plasmon mode with wave vector . We directly trace how the plasmon mode evolves from an unscreened plasmon, Eq. (1), for narrow gates to a screened plasmon for wider gates. Importantly, we have developed an analytical model that accurately describes the dispersion and relaxation of plasma waves propagating through the plasmonic crystal. Our results hold promise for engineering of THz plasmonic devices.
@@
Figure 2 shows the transmission through a series of samples with varying gate widths, keeping the grid period constant at µm. We refer to the Supplemental Material section for the analogous set of curves recorded for the plasmonic crystal with a period of µm [27]. For clarity, the curves are shifted along the vertical axis. Each sample exhibits several plasma resonances in the transmission spectrum. The dotted line marks the frequency position of the fundamental plasma resonance. As the gate width decreases, the frequency of the fundamental plasma mode increases. Figure 3 summarizes the frequency positions of the fundamental plasma mode versus the grid filling factor measured for the plasmonic crystal with µm (a) and µm (b). The dashed blue line in Fig. 3 depicts the frequency of the screened plasmon with calculated using Eq. (2). In addition, the dashed orange line in Fig. 3 depicts the frequency of the unscreened plasmon with from Eq. (1). In computations, we take due to the nonparabolicity of the electron conduction band occurring at such high 2D electron densities of [28, 29, 30]. It is obvious that the existing theoretical model marked by the blue and orange dashed lines [21, 22, 31, 23, 24] fails to describe the experimental results adequately.
It is natural to presume that at large , the plasmon should have a screened nature [10]. Indeed, Fig. 4(a) shows the spectrum of the plasma excitations observed for the sample with µm and period µm (). We now take the frequencies of the successive plasma peaks and assign them sequential wave vectors (). The resulting dispersion, displayed in Fig. 4(b) by red dots, bears a linear form characteristic of the screened plasmon. The theory (2) shown by a red line in Fig. 4(b) adequately agrees with the experimental points. In the opposite limit of narrow gates, when µm (), the dispersion obtained using the same procedure has a square-root dependency on (blue dots in Fig. 4(b)). Such behavior is characteristic for the unscreened plasmon with dispersion given by Eq. (1) and is presented by a blue line in Fig. 4(b). In the most common case of intermediate gate widths, the fundamental plasma state results from the hybridization of plasmon modes in gated and ungated parts of the 2DES.
In order to find the analytical expression for the frequency of the fundamental plasmon resonance, we introduce a standard method based on the solution of Maxwell’s equations for the plasmon electric and magnetic fields as well as the density currents arising in the system. The detailed derivation of the dispersion equation below can be found in the Supplemental Material [27].
When describing the plasmons, we assume that the gates are ideal metals, and we assume that the currents vanish at the edges. Also, we take into account that the wavelength of the THz radiation is much larger than the period of the plasmonic crystal, which greatly simplifies the analysis. Finally, considering that the plasmonic decay rate due to collisions with impurities and phonons as well as due to the emission of electromagnetic waves is small, one can find the following equation for the fundamental plasmon frequency :
@@
and can be expressed through special functions. Next, we left a sufficiently large number of terms (one hundred) in the sum in Eq. (3), after which the desired frequency can be calculated numerically.
@@
The frequencies of the fundamental plasma mode found from Eq. (3), are plotted as solid curves in Fig. 3 for different filling factors. There is a remarkable agreement between our theoretical predictions and the experimental results, which supports the validity of the approach to understand the fundamental plasma state in a plasmonic crystal. Several important observations can be made. It is particularly interesting that the plasmon frequency remains essentially constant, starting from the fully screened case of and all the way down to . In this state, the frequency of the plasma resonance stays close to the frequency of the screened plasmon Eq. (2) with (dotted horizontal line in Fig. 3). Beyond this point, the plasmon frequency increases sharply and eventually saturates near the filling factor of , aligning with the frequency of the unscreened plasmon with , Eq. (1), and the dielectric permittivity for µm and for µm (arrows in Fig. 3). Notably, the calculated values of the effective dielectric permittivity are slightly larger than the average value of . This difference arises most probably due to the µm GaAs top-layer covering the 2DES [27].
@@
Another matter of interest is associated with the relaxation dynamics of plasmons in the plasmonic crystal. Figure 5 shows the dependence of the fundamental plasma mode linewidth, , on the filling factor. For small filling factors, , the resonance linewidth is primarily influenced by a single-particle scattering. Indeed, a linewidth of GHz corresponds to a relaxation time of ps, which agrees with the electron mobility obtained from the transport measurements. However, as the filling factor increases, we observe a significant broadening of the resonance, accompanied by an increase in amplitude. This suggests that the plasmon linewidth is influenced by an additional mechanism beyond just single-particle scattering. The solid line in Fig. 5 shows the result of the theoretical analysis. 2D plasmons on a homogeneous surface are non-radiative modes. Nevertheless, the gate grating used for the coupling process also couples electromagnetic radiation out of the plasmon excitation [16, 3]. Such radiative damping is typical for plasmon excitations when the plasmon wavelength is comparable to the wavelength of the incident electromagnetic wave [32, 33]. However, the observation of radiative effects for the plasmon excitations in the limit is quite surprising.
@@
This expression provides an estimate of the maximum radiative decay rate that we can expect for our plasmonic crystal, GHz. The experimental data shown in Fig. 5 provides the radiative damping of GHz for the filling factor of . Therefore, the gate grating performs as an efficient radiative antenna.
In conclusion, we have investigated the plasma excitations in the terahertz transmission of the plasmonic crystal made from 2DES in the AlGaAs/GaAs semiconductor heterostructure with periodic metal grating deposited on the surface. Experiments have demonstrated that the plasmonic crystal supports only a single fundamental plasma mode with wave vector , which evolves from an unscreened state in a narrow-gate configuration to a screened state for wide gates. We find that the linewidth of the plasmon is determined by both the single-particle scattering and the radiative damping. An analytical approach has been developed which managed to accurately describe the fundamental plasma state in the plasmonic crystal. The obtained results establish a scientific foundation for engineering new plasmonic terahertz devices.
@@ Acknowledgements.
- [27] See supplemental material at [url will be inserted by publisher] for additional information about xxx, .

### v2

@@ Abstract
We report on the experimental study of the spectrum of plasma excitations in a plasmonic crystal fabricated from the two-dimensional electron system in an AlGaAs/GaAs semiconductor heterostructure. We perform a comprehensive research on the mode frequency and relaxation as a function of the gate width across different plasmonic crystal periods. Importantly, we develop an analytical approach that accurately describes the behavior of plasma excitations in plasmonic crystals, providing new insights into the fundamental physics of plasmonic systems.
@@
The currently prevailing theoretical model suggests that an incident electromagnetic wave excites two types of plasma waves in the plasmonic crystal. When the distance from the gate to the 2DES is relatively small, this model is analogous to the tight-binding limit of the Kronig-Penney model, which describes the particle behavior in a periodic potential [21, 22, 23, 24, 25]. The plasma wave in the ungated region, which is excited in the gap between the gates, is described by the spectrum [26]:
where is the 2D electron density, is the effective electron mass, is the unscreened plasmon wave vector, is the effective dielectric permittivity of the medium surrounding the 2DES. In the region directly underneath the gate, the plasma wave is the screened plasmon with the linear spectrum [27]
Here is the distance between the gate and the 2DES, is the screened plasmon wave vector, and is the dielectric permittivity of the semiconductor crystal, specifically . Then, the existing theoretical models make several assumptions. The plasmon dispersion relation for the system is obtained by (i) plane-wave expansion of the plasma wave electric potential in each region, (ii) matching of the potential and electric current at the gated-ungated boundaries, and (iii) imposing Bloch periodicity conditions for the fields in the neighboring gratings. These assumptions lead to Kronig-Penney-type dispersion of the optically active modes [21, 22, 23, 24, 25]:
where is the gate strips width, and is the width of the ungated 2DES area between the strips.
In the present work, we investigate the terahertz (THz) spectrum and relaxation of plasma excitations in a plasmonic crystal fabricated from an AlGaAs/GaAs heterostructure. Our experimental data reveal that the plasmonic crystal supports a series of resonant plasmon modes. We directly trace how the fundamental plasmon mode evolves from an unscreened plasmon, Eq. (1), for narrow gates to a screened plasmon, Eq. (2), for wider gates. Importantly, we have developed an analytical model that accurately describes the dispersion and relaxation of plasma waves propagating through the plasmonic crystal. The theory convincingly establishes that plasma excitations excited in the plasmonic crystal possess a delocalized superlattice nature. Our results hold promise for engineering of THz plasmonic devices.
@@
Figure 2 shows the transmission through a series of samples with varying gate widths, keeping the grid period constant at µm. We refer to the Supplemental Material section for the analogous set of curves recorded for the plasmonic crystal with a period of µm [28]. For clarity, the curves are shifted along the vertical axis. Each sample exhibits several plasma resonances in the transmission spectrum. The dotted line marks the frequency position of the fundamental plasma resonance, , as well as harmonic modes. As the gate width decreases, the frequency of the fundamental plasma mode increases. Figure 3 summarizes the frequency positions of the fundamental plasma mode versus the grid filling factor measured for the plasmonic crystal with µm (a) and µm (b). The dashed black lines in Fig. 3 depict the frequency of the plasmon modes predicted from the Kronig-Penney model Eq. (3) [28]. In computations, we take due to the nonparabolicity of the electron conduction band occurring at such high 2D electron densities of [29, 30, 31]. The results from model (3) show a reasonable agreement with the experimental data. However, since the model (3) is phenomenological in nature, the accuracy of the found resonant frequencies and ways for improvement are unclear. In addition, the model does not not account for the linewidth of the plasmon resonances, including the radiative contribution, nor does it provide information on the final line shape.
In order to find a better way to describe the behavior of plasmon modes in the plasmonic crystal, we develop a method based on the solution of Maxwell’s equations for space harmonics of the electric and magnetic fields:
where is the -th space harmonic of the grating, -axis is oriented perpendicular to the 2DES, -axis is oriented along the grating, and -axis is perpendicular to the grating. The detailed derivation of the dispersion equation below can be found in the Supplemental Material [28]. When describing the plasmons, we consider the gates to be ideal metals and assume that currents are negligible at the edges. Additionally, we note that the wavelength of the THz radiation is significantly larger than the period of the plasmonic crystal, which substantially simplifies our analysis. Lastly, given that the plasmon decay rate, caused by collisions with impurities, phonons, and the emission of electromagnetic waves, is minimal, we can derive the following equation for the fundamental plasmon frequency, :
@@
and can be expressed through the special functions. It is important to note that each plasmon mode of the plasmonic crystal consists of a superposition of standing plasma waves with wave-vectors propagating through the plasmonic crystal. To achieve good convergence, we left a sufficiently large number of space harmonics (one hundred) in the sum of Eq. (5), after which the desired plasmon frequency can be calculated numerically. The resulting frequencies of the fundamental plasmon mode, calculated from Eq. (5), are shown as solid curves in Fig. 3 for different filling factors. There is a remarkable agreement between theoretical predictions and experimental results, which supports the validity of our theoretical approach.
@@
Several important observations can be made. It is particularly interesting that the plasmon frequency remains essentially constant, starting from the fully screened case of and all the way down to . In this state, the frequency of the plasmon resonance stays close to the frequency of the screened plasmon Eq. (2) with (the empty arrows in Fig. 3). Beyond this point, the plasmon frequency increases sharply and eventually saturates near the filling factor of , aligning with the frequency of the unscreened plasmon with , Eq. (1), and the dielectric permittivity for µm and for µm (arrows in Fig. 3). Notably, the calculated values of the effective dielectric permittivity are slightly larger than the average value of . This difference arises due to the µm GaAs top-layer covering the 2DES [28].
To further illustrate the superlattice nature of the plasmon modes excited in the plasmonic crystal, we plot the normalized intensity of the calculated -th space harmonic of the fundamental plasmon mode for , and (inset of Fig. 4). The normalization is done with respect to the intensity of the fundamental harmonic. For small () and large () filling factors, the fundamental superlattice harmonic dominates the plasmonic response. Whereas, for intermediate -values, the and harmonics become significant. Indeed, Fig. 4(a) shows the spectrum of the plasma excitations observed for the sample with µm and period µm (). We now take the frequencies of the successive plasma peaks and assign them sequential wave vectors (). Here, we assume that only space harmonic contributes to the plasmonic response. The theoretical prediction, shown as a red line in Fig. 4(b), exhibits a linear form characteristic of the screened plasmon. The theory (2) shown by a red line in Fig. 4(b) adequately agrees with the experimental points. In the opposite limit of narrow gates, when µm (), the dispersion obtained using the same procedure has a square-root dependency on (blue dots in Fig. 4(b)). This behavior is typical of the unscreened plasmon, with its dispersion described by Eq. (1), represented by a blue line in Fig. 4(b). In the most common case of intermediate gate widths, the fundamental plasmon state arises from the hybridization of different superlattice space harmonics. Thus, a simple expression for the plasmon frequency cannot be provided.
@@
Another matter of interest is associated with the relaxation dynamics of plasmons in the plasmonic crystal. Figure 5 shows the dependence of the fundamental plasma mode linewidth, , on the filling factor. For small filling factors , the resonance linewidth is primarily influenced by a single-particle scattering. Indeed, a linewidth of GHz corresponds to a relaxation time of ps, which agrees with the electron mobility obtained from the transport measurements. However, as the filling factor increases, we observe a significant broadening of the resonance, accompanied by an increase in amplitude. This suggests that the plasmon linewidth is influenced by an additional mechanism beyond just single-particle scattering. The solid line in Fig. 5 shows the result of the theoretical analysis. 2D plasmons on a homogeneous surface are non-radiative modes. Nevertheless, the gate grating used for the coupling process also couples electromagnetic radiation out of the plasmon excitation [16, 3]. Such radiative damping is typical for plasmon excitations when the plasmon wavelength is comparable to the wavelength of the incident electromagnetic wave [32, 33]. However, the observation of radiative effects for the plasmon excitations in the limit is quite surprising.
@@
This expression provides an estimate of the maximum radiative decay rate that we can expect for our plasmonic crystal, GHz. As illustrated Fig. 5, the experimental data demonstrate an increase in radiative damping with an increase in filling factor . The range of values extends from GHz () to a maximum of GHz for a filling factor of . Therefore, the gate grating performs as an efficient radiative antenna.
In conclusion, we have investigated the plasma excitations in the terahertz transmission of the plasmonic crystal made from 2DES in the AlGaAs/GaAs semiconductor heterostructure with a periodic metal grating deposited on the surface. Experiments have demonstrated that the plasmonic crystal supports a series of plasmon modes, which evolve from an unscreened state in a narrow-gate configuration to a screened state for wide gates. An analytical approach has been developed that has managed to accurately describe the fundamental plasma state in the plasmonic crystal. It has been established that plasmons in a plasmonic crystal have a delocalized superlattice nature, being the combination of both screened and unscreened plasmons. Moreover, they consist of a superposition of multiple harmonics with wave vectors (). The obtained results establish a scientific foundation for engineering new plasmonic terahertz devices.
@@ Acknowledgements.
- [28] See supplemental material at [url will be inserted by publisher] for additional information about the samples, the experimantal data, and the theoretical model, .